OpenStudy (anonymous):

sqrt(18)a^2 A.2asqrt(2) B.(2)sqrt(9)a C.3asqrt(2) D.3asqrt(2)a whats the answer

OpenStudy (anonymous):

\[\sqrt{18a^2}?\]

OpenStudy (saifoo.khan):

\[\sqrt{18a^2}\]

OpenStudy (saifoo.khan):

is it like this?

OpenStudy (anonymous):

it's (C)

OpenStudy (anonymous):

if it is then 18=(3^2)(2) so the answer should b \[3a \sqrt{2}\]

OpenStudy (anonymous):

or C as that guy said